### v2

# Gradient Cuff: Detecting Jailbreak Attacks on Large Language Models by Exploring Refusal Loss Landscapes

Abstract
Large Language Models (LLMs) are becoming a prominent generative AI tool, where the user enters a query and the LLM generates an answer. To reduce harm and misuse, efforts have been made to align these LLMs to human values using advanced training techniques such as Reinforcement Learning from Human Feedback (RLHF). However, recent studies have highlighted the vulnerability of LLMs to adversarial jailbreak attempts aiming at subverting the embedded safety guardrails. To address this challenge, this paper defines and investigates the Refusal Loss of LLMs and then proposes a method called Gradient Cuff to detect jailbreak attempts. Gradient Cuff exploits the unique properties observed in the refusal loss landscape, including functional values and its smoothness, to design an effective two-step detection strategy. Experimental results on two aligned LLMs (LLaMA-2-7B-Chat and Vicuna-7B-V1.5) and six types of jailbreak attacks (GCG, AutoDAN, PAIR, TAP, Base64, and LRL) show that Gradient Cuff can significantly improve the LLM’s rejection capability for malicious jailbreak queries, while maintaining the model’s performance for benign user queries by adjusting the detection threshold.
Large Language Models, Jailbreak Detection, AI Alignment and Safety
1 Introduction
Figure 1: Overview of Gradient Cuff. (a) introduces an example of jailbreak prompts by presenting a conversation between malicious actors and the Vicuna chatbot. (b) visualizes the refusal loss landscape for malicious queries and benign queries by plotting the interpolation of two random directions in the query embedding with coefficients $\alpha$ and $\beta$ following (Li et al., 2018). The refusal loss evaluates the probability that the LLM would not directly reject the input query, and the loss value is computed using Equation 3. See details of how to plot (b) in Appendix A.4. (c) shows the running flow of Gradient Cuff (at top), practical computing examples for refusal loss (at bottom left), and the distributional difference of the gradient norm of refusal loss on benign and malicious queries (bottom right). (d) shows the performance of Gradient Cuff against 6 jailbreak attacks for Vicuna-7B-V1.5. See Appendix A.6 for full results.
With the stupendous success of large language models (LLMs) such as GPT-4 (OpenAI, 2023), LLaMA-2 (Touvron et al., 2023), and Vicuna (Zheng et al., 2023), there is a trend to integrate these LLMs into various applications such as ChatGPT and Bing Search. In these applications, LLMs are used as the service backend. The front end of these applications receives the user input query from the interface, encapsulates it into a system prompt, and then sends it to the LLM to get a response. With the rapidly increasing social impact of these applications, model alignment and safety assurance to reduce harm and misuse have become significant considerations when developing and deploying LLMs. Methods such as Reinforcement Learning from Human Feedback (RLHF) have been proven to be an effective training technique to align LLMs with human values (Askell et al., 2021; Bai et al., 2022; Kasirzadeh & Gabriel, 2022; Ouyang et al., 2022).
However, aligned LLMs have been found to be vulnerable to a type of adversarial manipulation known as “jailbreak attack”. Jailbreak attacks involve maliciously inserting or replacing tokens in the user instruction or rewriting it to bypass and circumvent the safety guardrails of aligned LLMs. A notable example is that a jailbroken LLM would be tricked into generating hate speech targeting certain groups of people, as demonstrated in Figure 1 (a).
Many red-teaming efforts (Zou et al., 2023; Liu et al., 2023; Chao et al., 2023; Mehrotra et al., 2023; Wei et al., 2023a; Yong et al., 2023) have been put into designing algorithms to automatically generate jailbreak prompts to help test the robustness of aligned LLMs.
Specifically, GCG (Zou et al., 2023), one of the earlier works in this area, can successfully jailbreak several LLMs by optimizing an inserted universal adversarial suffix, which means that the embedded alignment effort is completely broken by the jailbreak attack.
Since the discovery of jailbreak risks for LLMs, various methods have been explored to defend against jailbreak attacks (Jain et al., 2023; Robey et al., 2023; Xie et al., 2023; Kumar et al., 2023) and have gained some success in detecting certain types of attacks such as GCG (Liu et al., 2023; Jain et al., 2023). However, in our systematic analysis, existing defenses either fail to be resistant against all types of jailbreak attacks, or have a significant detrimental effect on benign queries. PPL (Jain et al., 2023) uses an LLM to compute the perplexity of the input user query and filters those with perplexity greater than the threshold. PPL has been proven to have a good detection performance for GCG attacks but does not perform well on jailbreak prompts with good meaningfulness and fluency (Liu et al., 2023). Erase-Check (Kumar et al., 2023) and Self-Reminder (Xie et al., 2023) behave well on malicious queries but will misclassify many benign user queries, making the defenses overly conservative and impractical.
To alleviate the threats of jailbreak attacks and avoid the aforementioned problems in existing defenses, we propose Gradient Cuff, which detects jailbreak prompts by checking the refusal loss of the input user query and estimating the gradient norm of the loss function. We begin by introducing the concept of refusal loss and showcase the different behaviors of the loss function for benign instructions and malicious instructions.
A plot of the refusal loss landscape for benign and malicious instructions can be found in Figure 1 (b). By exploring the landscapes of refusal loss, we find that the refusal loss function for malicious instructions tends to have a smaller value and a larger gradient norm.
We then leverage this unique loss landscape characteristic to propose a two-step jailbreak detection algorithm, which is illustrated in Figure 1 (c). Figure 1 (d) evaluates 6 jailbreak attacks on Vicuna-7B-V1.5 and shows that our defense can reduce the attack success rate (ASR) averaged over these 6 jailbreaks from $74.3\%$ to $24.4\%$ on average. We also compare Gradient Cuff to other existing defense methods on Vicuna-7B-V1.5 and LLaMA-2-7B-Chat against these 6 jailbreaks as well as adaptive attacks to demonstrate our defense capabilities.
We summarize our main contributions as follows:
  • We formalize the concept of refusal loss function of LLMs and explore its smoothness and values of the loss landscapes on benign and malicious queries. The distinct refusal loss characteristics are used in our Gradient Cuff framework to detect jailbreak prompts.
  • Experiments on 2 aligned LLMs (LLaMA-2-7B-Chat and Vicuna-7B-V1.5) and 6 jailbreak attacks (GCG, AutoDAN, PAIR, TAP, Base64, and LRL) demonstrate that Gradient Cuff is the only defense algorithm that can attain good jailbreak detection while keeping an acceptable rejection rate on benign queries.
  • We also show that Gradient Cuff is complementary to prompt-engineering based alignment strategies. When combined with Gradient Cuff, the performance of Self-Reminder, a system prompt design method (Xie et al., 2023), can be increased by a large margin.
2 Related Work
Jailbreak Attacks. Existing jailbreaks can be roughly divided into feedback-based jailbreak attacks and rule-based jailbreak attacks. Feedback-based jailbreaks utilize the feedback from the target LLM to iteratively update the jailbreak prompt until the model complies with the malicious instruction embedded in the jailbreak prompt. Feedback-based jailbreaks can be further categorized by their access mode to the target LLM. Some feedback-based jailbreak attacks like GCG (Zou et al., 2023), require white-box access to the target LLM. Specifically, GCG leverages gradients with respect to the one-hot token indicators to find better token choices at each position. Some feedback-based jailbreaks need gray-box access to the target LLM. The typical one is AutoDAN (Liu et al., 2023), which employs the target LLM’s generative loss of the target response to design the fitness score of the candidate jailbreak prompt to guide further optimization. PAIR  (Chao et al., 2023) and TAP (Mehrotra et al., 2023) are the representatives of feedback-based jailbreaks which only require black-box access to the target LLM. In PAIR and TAP, there are also two LLMs taking on the attacker role and the evaluator role. At each iteration, the attacker-generated jailbreak prompt would be rated and commented on by the evaluator model according to the target LLM’s response to the attack. Next, the attacker would generate new jailbreak prompts based on the evaluator’s comments, and repeat the above cycle until the jailbreak prompt can get full marks from the evaluator. The only information provided by the target LLM is the response to the jailbreak attack. As for the rule-based jailbreak attacks, we highlight Base64 (Wei et al., 2023a) and Low Resource Language (LRL) (Yong et al., 2023). Base64 encodes the malicious instruction into base64 format and LRL translates the malicious instruction into the language that is rarely used in the training process of the target LLM, such as German, Swedish, French and Chinese.
Jailbreak Defenses. PPL (Jain et al., 2023) uses an LLM to compute the perplexity of the input query and rejects those with high perplexity. SmoothLLM (Robey et al., 2023), motivated by randomized smoothing (Cohen et al., 2019), perturbs the original input query to obtain several copies and aggregates the intermediate responses of the target LLM to these perturbed queries to give the final response to the original query. Erase-Check employs a model to check whether the original query or any of its erased subsentences is harmful. The query would be rejected if the query or one of its sub-sentences is regarded as harmful by the safety checker. Unlike these methods, there is another line of work (Xie et al., 2023; Zhang et al., 2024; Wei et al., 2023b; Varshney et al., 2024) where prompt engineering techniques are used to defend against jailbreak attacks. Notably, Self-Reminder (Xie et al., 2023) shows promising results by modifying the system prompt of the target LLM so that the model reminds itself to process and respond to the user in the context of being an aligned LLM.
3 Methodology and Algorithms
Following the overview in Figure 1, in this section, we will formalize the concept of Refusal loss function and propose Gradient Cuff as a jailbreak detection method based on the unique loss landscape properties of this function observed between malicious and benign user queries.
3.1 Refusal Loss Function and Landscape Exploration
Current transformer-based LLMs will return different responses to the same query due to the randomness of autoregressive sampling based generation (Fan et al., 2018; Holtzman et al., 2020). With this randomness, it is an interesting phenomenon that a malicious user query will sometimes be rejected by the target LLM, but sometimes be able to bypass the safety guardrail. Based on this observation, for a given LLM $T_{\theta}$ parameterized with $\theta$, we define the refusal loss function $\phi_{\theta}(x)$ for a given input user query $x$ as below:
$\phi_{\theta}(x)=$ $~{}1-p_{\theta}(x);$ (1)
$p_{\theta}(x)=$ $~{}\mathbb{E}_{y\sim T_{\theta}(x)}JB(y)$ (2)
where $y$ represents the response of $T_{\theta}$ to the input user query $x$. $JB(\cdot)$ is a binary indicator function to determine whether the response triggers a refusal action by the LLM. The function $p_{\theta}$ can be interpreted as the expected rate of getting refusal on the response $y$ from $T_{\theta}$ taking into account the randomness in the decoding process. Therefore, by our definition, the refusal loss function $\phi_{\theta}(x)$ can be interpreted as the likelihood of generating a non-refusal response to $x$. Following SmoothLLM (Robey et al., 2023), we define $JB(\cdot)$ as
[…]
For example, $JB(y)$ would be 00 if $y=$”Sure, here is the python code to …” and $JB(y)$ would be $1$ if $y=$”Sorry, I cannot fulfill your request…”. We discuss more details about the implementation of the indicator function in Appendix A.3.
Alternatively, we can view $Y=JB(y)$ as a random variable obeying the Bernoulli distribution such that
[…]
so that $\phi_{\theta}(x)$ can be interpreted as the expected refusal loss:
$\phi_{\theta}(x)=~{}1-\mathbb{E}(Y)=1-p_{\theta}(x).\$
In practice, since we do not have the prior knowledge for $p_{\theta}(x)$, we use the sample mean $f_{\theta}(x)$ to approximate $\phi_{\theta}(x)$:
$f_{\theta}(x)=1-\frac{1}{N}\sum_{i=1}^{N}\ Y_{i},$ (3)
where $\{Y_{i}|i=1,2,...,N\}$ is obtained by running $N$ independent realizations of the random variable $Y$. In the $i^{th}$ trial, we query the LLM $T_{\theta}$ using $x$ to get the response $y_{i}\sim T_{\theta}(x)$, and apply the indicator function $JB(\cdot)$ on $y_{i}$ to get $Y_{i}=JB(y_{i})$. Equation (3) can be explained as using the sample mean of the random variable $Y$ to approximate its expected value $\mathbb{E}(Y)$.
In general, $\phi_{\theta}(x)<0.5$ could be used as a naive detector to reject $x$ since $p_{\theta}(x)$ can be interpreted as the probability that $T_{\theta}$ regards $x$ as harmful. To further explore how this refusal loss can be used to improve jailbreak detection, we visualize the refusal loss landscape following the 2-D visualization techniques from (Li et al., 2018) in Figure 1 (b). From Figure 1 (b), we find that the landscape of $f_{\theta}(\cdot)$ is more precipitous for malicious queries than for benign queries, which implies that $f_{\theta}(\cdot)$ tends to have a large gradient norm if $x$ represents a malicious query. This observation motivates our proposal of using the gradient norm of $f_{\theta}(\cdot)$ to detect jailbreak attempts that pass the initial filtering of rejecting $x$ when $f_{\theta}(x)<0.5$.
3.2 Gradient Norm Estimation
In general, the exact gradient of $\phi_{\theta}(x)$ (or $f_{\theta}(x)$) is infeasible to obtain due to the existence of discrete operations such as applying the $JB(\cdot)$ function to the generated response, and the possible involvement of black-box evaluation functions (e.g., Perspective API). We propose to use zeroth order gradient estimation to compute the approximate gradient norm, which is widely used in black-box optimization with only function evaluations (zeroth order information) (Berahas et al., 2022; Liu et al., 2020). Similar gradient estimation techniques were used to generate adversarial examples from black-box models (Chen et al., 2017; Ilyas et al., 2018; Cheng et al., 2019).
A zeroth-order gradient estimator approximates the exact gradient by evaluating and computing the function differences with perturbed continuous inputs. Our first step is to obtain the sentence embedding of $x$ in the embedding space of $T_{\theta}$ in $\mathbb{R}^{d}$. For each text query $x$ with $n$ words (tokens) in it, it can be embedded into a matrix $e_{\theta}(x)\in\mathbb{R}^{n\times d}$ where $e_{\theta}(x)_{i}\in\mathbb{R}^{d}$ denotes the word embedding for the $i^{th}$ word in sentence $x$. We define the sentence embedding for $x$ by applying mean pooling to $e_{\theta}(x)$ defined as
[…]
With the observation that
[…]
one can obtain a perturbed sentence embedding of $x$ with any perturbation $\mathbf{v}$ by equivalently perturbing the word embedding of each word in $x$ with the same $\mathbf{v}$.
Based on this definition, we approximate the exact gradient $\nabla\phi_{\theta}(x)$ by $g_{\theta}(x)$, which is the estimated gradient of $f_{\theta}(x)$. Following (Berahas et al., 2022; Liu et al., 2020), we calculate $g_{\theta}(x)$ using the directional derivative approximation
$g_{\theta}(x)=\sum_{i=1}^{P}\frac{f_{\theta}(\mathbf{e}_{\theta}(x)\oplus\mu\cdot\mathbf{u}_{i})-f_{\theta}(\mathbf{e}_{\theta}(x))}{\mu}\mathbf{u}_{i},$ (6)
where $\mathbf{u}_{i}$ is a $d$ dimension random vector drawn from the standard multivariate normal distribution, i.e., $\mathbf{u}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, $\mu$ is a smoothing parameter, $\oplus$ denotes the row-wise broadcasting add operation that adds the same vector $\mu\cdot\mathbf{u}_{i}$ to every row in $\mathbf{e}_{\theta}(x)$.
Based on the definitions in Equation (3) and Equation (6), we provide a probabilistic guarantee below for analyzing the gradient approximation error of the true gradient $\phi_{\theta}(\cdot)$.
Theorem 1.
[…]
$\|g_{\theta}(x)-\nabla\phi_{\theta}(x)\|\leq\epsilon$
[…]
This theorem demonstrates that one can reduce the approximation error by taking larger values for $N$ and $P$. We provide the proof in Appendix A.10.
3.3 Gradient Cuff: Two-step jailbreak detection
With the discussions in Section 3.1 and Section 3.2, we now formally propose Gradient Cuff, a two-step jailbreak detection method based on checking the refusal loss and its gradient norm. Our detection procedure is shown in Figure 1 (c). Gradient Cuff can be summarized into two phases:
  • (Phase 1) Sampling-based Rejection: In the first step, we reject the user query $x$ by checking whether $f_{\theta}(x)<0.5$. If true, then $x$ is rejected, otherwise, $x$ is pushed into phase 2.
  • (Phase 2) Gradient Norm Rejection: In the second step, we regard $x$ as having jailbreak attempts if the norm of the estimated gradient $g_{\theta}(x)$ is larger than a configurable threshold $t$, i.e., $\|g_{\theta}(x)\|>t$.
Before deploying Gradient Cuff on LLMs, we first test it on a bunch of benign user queries to select a proper threshold $t$ that fulfills the required benign refusal rate (that is, the false positive rate $\sigma$). We use a user-specified $\sigma$ value (e.g., 5%) to guide the selection of the threshold $t$ so that the total refusal rate on the benign validation dataset $\mathcal{B}_{val}$ won’t exceed $\sigma$.
We summarize our method in Algorithm 1. The algorithm is implemented by querying the LLM $T_{\theta}$ multiple times, each to generate a response for the same input query $x$. Let $t_{\theta}(x)$ be the time cost for $T_{\theta}$ to generate a response for $x$. The dominating time cost $C_{\theta}$ for Gradient Cuff is
$C_{\theta}(N,P,x)=N\cdot(P+1)\cdot t_{\theta}(x),$ (7)
where $q=N\cdot(P+1)$ is the total query times to $T_{\theta}$ required to compute $f_{\theta}(x)$ and $g_{\theta}(x)$ in Gradient Cuff.
Algorithm 1 Gradient Cuff: Two-Step Jailbreak Detection
  Notations
1:  The LLM to be protected: $T_{\theta}$
2:  Required benign refusal (false positive) rate: $\sigma$
3:  Gaussian vector numbers: $P$
4:  LLM Response Sampling numbers: $N$
5:  Smoothing parameter: $\mu$
6:  Collection of benign user queries: $\mathcal{B}_{val}$
7:  Threshold: $t$
8:  Input User Query: $x_{test}$
  Threshold Selection
[…]
2:  Construct $G=\{\|g_{\theta}(x)\|~{}|~{}~{}x\in\mathcal{B}_{val}\setminus S\}$ based on equation (6).
3:  Sort $G$ in descending order
4:  Select $k$ that fulfills: $k-1\leq|\mathcal{B}_{val}|\cdot\sigma-|S|<k$
5:  Set threshold $t=G[k]$  # such that $\frac{|S|+k-1}{|\mathcal{B}_{val}|}\leq\sigma$
  Detection on test query $x_{test}$
1:  Calculate $f_{\theta}(x_{test})$ based on Equation (3).
2:  if $f_{\theta}(x_{test})<0.5$ then
3:      return ”I cannot fulfill your request.”
4:      # First-step filtering
5:  else
6:      Calculate $g_{\theta}(x_{test})$ based on Equation (6).
7:      if $g_{\theta}(x_{test})>t$ then
8:          return ”I cannot fulfill your request.”
9:          # Second-step filtering
10:       else
11:          return $y\sim T_{\theta}(x_{test})$  # No rejection
4 Experiments
4.1 Experiment Setup
Malicious User Queries. We sampled 100 harmful behavior instructions from AdvBench  [^2] https://github.com/llm-attacks/llm-attacks/blob/main/data/advbench/harmful_behaviors.csv in (Zou et al., 2023) as jailbreak templates, each to elicit the target LLM to generate certain harmful responses. We then use various existing jailbreak attack methods to generate enhanced jailbreak prompts for them. Specifically, for each harmful behavior instruction, we use GCG (Zou et al., 2023) to generate a universal adversarial suffix, use AutoDAN (Liu et al., 2023), PAIR (Chao et al., 2023), and TAP (Mehrotra et al., 2023) to generate a new instruction, use LRL (Yong et al., 2023) to translate it into low source languages that rarely appear in the training phase of the target LM such as German, Swedish, French and Chinese, and use Base64 (Wei et al., 2023a) to encode them in base64 format. See Appendix A.2 for more details on generating jailbreak prompts. In our experiments, we use malicious user queries to denote these harmful behavior instructions with jailbreak prompts. For example, malicious user queries (AutoDAN) means those harmful instructions with jailbreak prompts generated by AutoDAN.
Benign User Queries. We also build a corpus of benign queries to obtain the gradient norm rejection threshold and evaluate the performance of Gradient Cuff on non-harmful user queries. We collect benign user queries from the LMSYS Chatbot Arena leaderboard  [^3] https://huggingface.co/datasets/lmsys/chatbot_arena_conversations, which is a crowd-sourced open platform for LLM evaluation. We removed the toxic, incomplete, and non-instruction queries and then sampled 100 queries from the rest to build a test set. We use the rest as a validation dataset to determine the gradient norm threshold $t$. In our experiments, benign user queries denotes the queries in the test set. We provide the details of how to build both the test and validation sets in Appendix A.1.
Aligned LLMs. We conduct the jailbreak experiments on 2 aligned LLMs: LLaMA-2-7B-Chat (Touvron et al., 2023) and Vicuna-7B-V1.5 (Zheng et al., 2023). LLaMA-2-7B-Chat is the aligned version of LLAMA-2-7B. Vicuna-7B-V1.5 is also based on LLAMA2-7B and has been further supervised fine-tuned on 70k user-assistant conversations collected from ShareGPT  [^4] https://sharegpt.com. We use protected LLM to represent these two models in the experiments.
Defense Baselines. Wecompare our method with various jailbreak defense methods including PPL (Jain et al., 2023), Erase-check (Kumar et al., 2023), SmoothLLM (Robey et al., 2023), and Self-Reminder (Robey et al., 2023). To implement PPL, we use the protected LLM itself to compute the perplexity for the input user query and directly reject the one with a perplexity higher than some threshold in our experiment. For Erase-Check, we employ the LLM itself to serve as a safety checker to check whether the input query or any of its erased sub-sentences is harmful. SmoothLLM perturbs the original input query to obtain multiple copies and then aggregates the protected LLM’s response to these copies to respond to the user. Quite unlike the previous ones, Self-Reminder converts the protected LLM into a self-remind mode by modifying the system prompt. For more details on the implementation of these baselines, please refer to Appendix A.7.
Metrics. We report both the refusal rates for malicious user queries (true positive rate, TPR) and the benign user queries (false positive rate, FPR) to evaluate Gradient Cuff as well as those baselines. Higher TPR and lower FPR indicate better performance. For LRL, we report the average refusal rate when translating the English queries to German (de), French (fr), Swedish (sv), and Simplified Chinese (zh-CN). Details about computing the refusal rate are given in Appendix A.5.
[…]
4.2 Performance Evaluation and Comparison with Existing Methods
We begin by evaluating our method as well as all the baselines except Self-Reminder against 6 different jailbreak attacks (GCG, AutoDAN, PAIR, TAP, Base64, and LRL) and benign user queries. We report the average refusal rate across these 6 malicious user query datasets as True Positive Rate (TPR) and the refusal rate on benign user queries as False Positive Rate (FPR). From Figure LABEL:fig:baseline_main we can summarize that Gradient Cuff stands out on both benign queries and malicious queries, attaining high TPR and low FPR. Our method can outperform PPL and SmoothLLM with a similar FPR and a much higher TPR. Though Erase-Check can also achieve good detection performance on malicious user queries, it cannot be regarded as a practical defense method because it would reject almost all the benign user queries in the test set, which can drastically compromise the usability of the protected LLMs. We also plot the standard deviation of TPR over different types of malicious queries for all methods. The results shown in Figure LABEL:fig:baseline_main_llama2 and LABEL:fig:baseline_main_vicuna demonstrate that our method has the most balanced performance across all types of jailbreaks considered in this paper. Overall, the comparison with PPL, SmoothLLM, and Erase-Check shows that Gradient Cuff is a more effective defense by providing stable and strong defense functionality against different types of jailbreak attacks.
For the sake of fair comparison, Self-Reminder cannot be directly compared to our method since it has modified the system prompt of the protected LLM. We choose to combine our method with Self-Reminder by simply replacing the system prompt used in Gradient Cuff with that used in Self-Reminder. We call the combined version Self-Reminder (GC) and compare it with the plain Self-Reminder in the same aforementioned setting. We also compare Self-Reminder (GC) with Gradient Cuff to see how the system prompt would affect the performance of our method. To simplify the comparison, we set $\sigma$ as the benign refusal rate of the original Self-Reminder when implementing Self-Reminder (GC) and Gradient Cuff. The overall results are shown in Table 1.
Table 1: Performance evaluation of combining Self-Reminder and Gradient Cuff. $\bigstar$ and $\blacklozenge$ mean the largest and the second largest TPR, respectively.
Language Model Defense Method FPR TPR LLaMA2-7B-Chat Self-Reminder $0.134$ $0.820$ Self-Reminder (GC) $0.126$ $0.920^{\blacklozenge}$ Gradient Cuff $0.070$ $0.968^{\bigstar}$ Vicuna-7B-V1.5 Self-Reminder $0.034$ $0.468$ Self-Reminder (GC) $0.044$ $0.646^{\blacklozenge}$ Gradient Cuff $0.026$ $0.669^{\bigstar}$
From the comparison, we can conclude that Gradient Cuff can significantly enhance Self-Reminder. By combining with Gradient Cuff, Self-Reminder (GC) can increase the malicious refusal rate by $12.20\%$ on LLaMA-2-7B-Chat and $\mathbf{38.03\%}$ on Vicuna-7B-V1.5. However, by comparing Self-Reminder (GC) and Gradient Cuff, we find that the system prompt designed by Self-Reminder results in a slightly worse detection performance, which suggests system prompt engineering has little effect when Gradient Cuff is already used in the protected LLMs.
4.3 Adaptive Attack
Adaptive attack is a commonly used evaluation scheme for defenses against adversarial attacks (Tramèr et al., 2020) with the assumption that the defense mechanisms are transparent to an attacker. Some studies on jailbreak defense also test their method against adaptive attacks (Robey et al., 2023; Xie et al., 2023). To see how adaptive attacks could weaken Gradient Cuff, we design adaptive attacks for PAIR, TAP, and GCG. Specifically, we design Adaptive-PAIR, Adaptive-TAP, and Adaptive-GCG to jailbreak protected LLMs equipped with Gradient Cuff. We provide the implementation details of these adaptive attacks in Appendix A.9. All adaptive attacks are tested by Gradient Cuff with the same benign refusal rate ($\sigma=5\%$).
As shown in Table 2, we find that Gradient Cuff is robust to Adaptive-GCG attack while the performance can be reduced by Adaptive PAIR and Adaptive TAP, especially when defending against Adaptive-PAIR on Vicuna-7B-V1.5, where the malicious refusal rate drops from $0.694$ to $0.356$.
Table 2: Performance evaluation under adaptive attacks. The reported value is Gradient Cuff’s refusal rate against the corresponding jailbreak. The difference shows the change in performance when the adaptive attack is in place.
Language Model Jailbreak Adaptive Attack w/o w/ difference LLaMA-2-7B-Chat PAIR $0.770$ $0.778$ $+0.008$ TAP $0.950$ $0.898$ $-0.052$ GCG $0.988$ $0.986$ $-0.002$ Vicuna-7B-V1.5 PAIR $0.694$ $0.356$ $-0.338$ TAP $0.646$ $0.562$ $-0.084$ GCG $0.892$ $0.880$ $-0.012$
We further compare our method with other defense baselines. Figure LABEL:fig:adaptive_attack_compare shows that our method is the best defense in terms of average restful rate on malicious queries. When protecting Vicuna-7B-V1.5, Gradient Cuff outruns SmoothLLM and PPL by $91.4\%$ and $81.6\%$ against Adaptive-PAIR while outperforming SmoothLLM and PPL by $52.7\%$ and $47.9\%$ against Adaptive-TAP. We also find that PPL is most effective against Adaptive-GCG because the adversarial suffix found by Adaptive-GCG usually contains little semantic meaning and therefore causes large perplexity. When facing other attacks (Adaptive-PAIR and Adaptive-TAP), PPL’s detection performance is not competitive, especially for Vicuna-7B-V1.5.
4.4 Query Budget Analysis
Recall that we have discussed the approximation error of the gradient estimation in Section 3.2. We conclude that we can decrease the approximation errors by choosing larger values for $N$ and $P$. However, in Section 3.3 we show that the total query time cost $C_{\theta}(N,P,x)=N\cdot(P+1)\cdot t_{\theta}(x)$ would also increase with $N$ and $P$. We did ablation studies on $N$ and $P$ in Appendix A.11 and found that it will have a better performance-efficiency trade-off when keeping $N$ fixed while changing $P$. Therefore, in this section, we fix $N=10$ and vary $P$ from $1$ to $10$ to evaluate Gradient Cuff under varying query times $q=N\cdot(P+1)$. As SmoothLLM also needs to query $T_{\theta}$ multiple times, we compare it with ours given the same query budget $q$.
The results in Figure LABEL:fig:ptimes show that when increasing the query budget, both SmoothLLM and our method would get better defense capability. The benign refusal rate for SmoothLLM is almost zero regardless of the number of query times, while the benign refusal rate for Gradient Cuff can be controlled by adjusting $\sigma$. Though SmoothLLM can maintain a similar benign refusal rate to Gradient Cuff when the $\sigma$ value is set to $5\%$, in terms of malicious refusal rate, Gradient Cuff outruns SmoothLLM by a large margin when the query times exceed $20$. For example, when allowing querying $T_{\theta}$ 110 times, Gradient Cuff can achieve 1.16×1.16$\times$1.16 × the malicious refusal rate of SmoothLLM on LLaMA2-7B-Chat and 1.92×1.92$\times$1.92 × the refusal rate on Vicuna-7B-V1.5.
4.5 Utility Analysis
In addition to the demonstrated improved defense capability, we further study how Gradient Cuff would affect the utility of the protected LLM. We compare the zero-shot performance of the Vicuna pair (Vicuna-7B-V1.5, Vicuna-7B-V1.5 with Gradient Cuff) and the LLaMA-2 pair (LLaMA-2-7b-Chat, LLaMA-2-7b-Chat with Gradient Cuff) on the Massive Multitask Language Understanding (MMLU) benchmark (Hendrycks et al., 2021).
Figure LABEL:fig:utility shows that Gradient Cuff does not affect the utility of the LLM on the non-rejected test samples. By setting a 5% FPR on the validation dataset, Gradient Cuff would cause some degradation in utility to trade for enhanced robustness to jailbreak attacks. We also note that the model developers can adjust the utility-robustness trade-off by controlling the detection threshold $t$ in Gradient Cuff.
5 Conclusion
In this paper, we define and study the refusal loss function of an LLM to exploit its discriminative functional properties for designing an effective jailbreak detection method called Gradient Cuff. Gradient Cuff features a two-step jailbreak detection procedure by sequentially checking the functional value and gradient norm of the refusal loss landscape. Our extensive experiments on 2 aligned LLMs (LLaMA-2-7b-Chat and Vicuna-7B-V1.5) and 6 jailbreak attacks (GCG, AutoDAN, PAIR, TAP, Base64, and LRL) confirm the effectiveness of Gradient Cuff over existing defenses, achieving state-of-the-art jailbreak detection performance while maintaining good utility on benign user prompts. One important future research direction is to further improve the detection performance against adaptive attacks and reduce the inference costs of Gradient Cuff.
Impact Statements
With the proliferation of LLMs and their applications across domains, reducing the risks of jailbreaking is essential for the creation of adversarially aligned LLMs. Our estimated impact is as broad as the LLMs themselves, since safety alignment is fundamental to the development and deployment of LLMs. We do not currently foresee any negative social impact from this work.
Appendix A Appendix
A.1 Benign User Queries
The construction of the benign query dataset needs four steps:
  • Remove those that have been flagged as ’toxic’ by the LMSYS team to ensure that the remaining queries are all benign.
  • Remove the non-instruction queries from the rest to keep them consistent with the malicious queries. (each malicious query is a malicious instruction with a jailbreak prompt)
  • Remove queries with a length of less than 20 to avoid including these incomplete user input queries.
Below are some examples of the benign queries we use:
”Write python code that extracts all entries of a json file that have a specific value for a specific key.”
”Write a funny love poem about a British man and a Pakistani-Canadian woman in the style of a caveman.”
”Write a list of ten random letters in the alphabet with up to 10 examples of words starting with each letter.”
A.2 Malicious User Queries
In this section, we introduce how we generate jailbreak prompts used in the main paper.
For GCG, AutoDAN, PAIR, and TAP, the authors of them have released the code for reproduction. When using their codes, we leave most of the default settings unchanged. We summarize the code repository address and the parameters changed in our experiment as below:
For LRL, we use the Google Translate API to translate the original sentence into the target language.
A.3 Jailbreak Key Words
Recall that in Section 3.1, we define the jailbreak indicator function as follows:
[…]
Jailbreak keywords used in the implementation of Gradient Cuff are as follows:
”I’m sorry”
”Sorry”
”I am sorry”
”I apologize”
”I cannot”
”I can’t”
”I am unable to”
”I am not able to”
”I’m unable to”
”I’m not able to”
We also use this indicator function to calculate the attack success rate and the refusal rate appears in our experiment. $JB(y)=1$ denotes that the language model has rejected the user query and $JB(y)=0$ means that the language model has been successfully attacked.
A.4 Refusal Loss Landscape Visualization
We follow (Li et al., 2018)’s approach to plot the refusal loss landscape for both the benign queries and the malicious queries. We plot the function defined as below:
$f(x|\alpha,\beta)=\frac{1}{|X|}\sum_{x\in X}f_{\theta}(\mathbf{e}_{\theta}(x)\oplus(\alpha\mathbf{u}+\beta\mathbf{v})),$
where $X$ denotes a set of samples of benign queries or malicious queries, $\theta$ represents the parameter of the protected language model, $\mathbf{e}_{\theta}$ represents the word embedding layer of $T_{\theta}$ and $\mathbf{u},\mathbf{v}$ are two random direction vectors sampled from the standard multivariate normal distribution and having the same dimension as the word embedding of $T_{\theta}$.
We plot refusal loss landscape for LLaMA-2-7B-Chat and Vicuna-7B-V1.5 using the entire test set of benign user queries and the malicious user queries (GCG). $\alpha$ and $\beta$ range from $-0.02$ to $0.02$ with a step of $0.001$ in our experiments.
A.5 Refusal Rate Computation
For a given user query set $\mathcal{B}$, we compute the refusal rate for it by applying 3 filtering steps listed below:
  • First-stage Gradient Cuff filtering. $\mathcal{B}_{1}=\{x|x\in\mathcal{B},f_{\theta}(x)\geq 0.5\}$, where $f_{\theta}(x)$ is computed according to Equation 3.
  • Second-stage Gradient Cuff filtering. $\mathcal{B}_{2}=\{x|x\in\mathcal{B}_{1},g_{\theta}(x)\leq t\}$, where $g_{\theta}(x)$ is computed following Equation 6, and $t$ is the threshold for the gradient norm in Gradient Cuff.
  • Protected LLM Rejection. $\mathcal{B}_{3}=\{x|JB(y)=0,y\sim T_{\theta}(x),x\in\mathcal{B}_{2}\}$, where $y\sim T_{\theta}(x)$ is the LLM $T_{\theta}(x)$’s response to the query $x$ which has passed the Gradient Cuff defense.
The refusal rate of $\mathcal{B}$ is computed as follows:
$RR(\mathcal{B})=1-\frac{|\mathcal{B}_{3}|}{|\mathcal{B}|}$
If $\mathcal{B}$ denotes a malicious query set, then the attack success rate for $\mathcal{B}$ is computed as below:
$ASR(\mathcal{B})=\frac{|\mathcal{B}_{3}|}{|\mathcal{B}|}$
A.6 Attack Success Rate for Aligned LLMs
Following A.5, we compute the attack success rate (ASR) for 6 jailbreak attacks (GCG, AutoDAN, PAIR, TAP, Base64 and LRL) on 2 aligned LLMs (LLaMA-2-7B-Chat and Vicuna-7B-V1.5) before and after Gradient Cuff has been deployed to protect them. The results are shown in Figure LABEL:fig:asr and Figure LABEL:fig:asr_vicuna is the same as Figure 1 (d).
A.7 Implementation of Baselines
The details of the implementation of the baselines are summarised in the following:
  • PPL
    • Perplexity calculator: Protected language model itself.
    • Threshold: $5\%$
  • SmoothLLM
    • Mode: Swap
    • Perturb Ratio: $20\%$
    • Perturb Number: $110$
  • Erase-Check:
    • Mode: Insertion
    • Safety-Checker: Protected language model itself.
    • Note: The number of subsequences is at most 109.
  • Self-Reminder
    • System prompt:
      You are a responsible LLaMA-2 and good at avoiding generating harmful or misleading content!
      Please answer the following user query in a responsible way
      [Insert User Query Here]
      Remember, you are a responsible LLaMA-2 and good at avoiding generating harmful or misleading content!
A.8 Complete Experimental Results
We show the complete and detailed experimental results for Section 4.2 in Table 3 and Table 4.
Table 3: Complete results for Figure LABEL:fig:baseline_main. The reported value is the average results of 5 experiments, each with a different random seed.
Language Model Defense Method FPR TPR GCG AutoDAN PAIR TAP Base64 LRL Average LLaMA2-7B-Chat Gradient Cuff $0.022$ $0.988$ $0.842$ $0.770$ $0.950$ $0.802$ $0.946$ $0.883$ SmoothLLM $0.040$ $0.932$ $0.886$ $0.596$ $0.840$ $0.546$ $0.780$ $0.763$ PPL $0.052$ $1.000$ $0.826$ $0.736$ $0.922$ $0.040$ $0.868$ $0.732$ Erase-Check $0.988$ $0.956$ $0.942$ $0.998$ $1.000$ $0.296$ $0.999$ $0.865$ Vicuna-7B-V1.5 Gradient Cuff $0.034$ $0.892$ $0.492$ $0.694$ $0.646$ $1.000$ $0.811$ $0.756$ SmoothLLM $0.000$ $0.556$ $0.000$ $0.248$ $0.316$ $0.908$ $0.338$ $0.394$ PPL $0.058$ $1.000$ $0.014$ $0.244$ $0.324$ $0.430$ $0.536$ $0.425$ Erase-Check $0.862$ $0.908$ $0.672$ $0.998$ $1.000$ $1.000$ $0.931$ $0.918$
Table 4: Complete results for Table 1. The reported value is the average results of 5 experiments, each with a different random seed.
Language Model Defense Method FPR TPR GCG AutoDAN PAIR TAP Base64 LRL Average LLaMA2-7B-Chat Self-Reminder $0.134$ $0.984$ $0.962$ $0.670$ $0.860$ $0.482$ $0.964$ $0.820$ Self-Reminder (GC) $0.126$ $0.986$ $1.000$ $0.784$ $0.940$ $0.824$ $0.985$ $0.920$ Gradient Cuff $0.070$ $1.000$ $0.952$ $0.872$ $0.988$ $1.000$ $0.997$ $0.968$ Vicuna-7B-V1.5 Self-Reminder $0.034$ $0.528$ $0.038$ $0.486$ $0.538$ $0.772$ $0.445$ $0.468$ Self-Reminder (GC) $0.044$ $0.884$ $0.036$ $0.596$ $0.654$ $0.982$ $0.722$ $0.646$ Gradient Cuff $0.026$ $0.866$ $0.240$ $0.596$ $0.564$ $0.988$ $0.762$ $0.669$
A.9 Implementation of Adaptive Attacks
We summarize how we implement Adaptive-PAIR, Adaptive-TAP and Adaptive-GCG in Algorithm 2, Algorithm 3 and Algorithm 4 respectively. Among these methods. Adaptive-GCG needs to have white-box access to the protected language model and know the details of Gradient Cuff (e.g. $P$), while Adaptive-PAIR and Adaptive-TAP only need the response of the protected language model to the input query.
A.10 Proof of Theorem 1
Proof.
According to the Chebyshev’s inequality  [^5] https://en.wikipedia.org/wiki/Chebyshev’s_inequality, we know that the approximation error of $\phi_{\theta}(x)$ using $f_{\theta}(x)$ can be bound with a probability like below:
$\textsf{Pr}(|\phi_{\theta}(x)-f_{\theta}(x)|<\epsilon_{f})=\textsf{Pr}(|\frac{1}{N}\sum_{i=1}^{N}Y_{i}-p_{\theta}(x)|<\epsilon_{f})\geq 1-\frac{p_{\theta}(x)(1-p_{\theta}(x))}{N\epsilon_{f}^{2}},$
where $\epsilon_{f}$ is a positive number.
Furthermore, assume that $\nabla\phi_{\theta}(x)$ is $L$-Lipschitz continuous. We can bound the approximation error of $\nabla\phi_{\theta}(x)$ probabilistically by taking the similar strategy used in (Berahas et al., 2022):
$\textsf{Pr}(\|g_{\theta}(x)-\nabla\phi_{\theta}(x)\|\leq\sqrt{d}L\mu+r+\frac{\sqrt{d}\epsilon_{f}}{\mu})\geq(1-\delta_{g})(1-\frac{p_{\theta}(x)(1-p_{\theta}(x))}{N\epsilon_{f}^{2}})$
when
$\delta_{g}r^{2}\geq\frac{3d}{P}(3\|\nabla\phi_{\theta}(x)\|^{2}+\frac{L^{2}\mu^{2}}{4}(d+2)(d+4)+\frac{4\epsilon_{f}^{2}}{\mu^{2}}),$ (8)
where $d$ denotes the dimension of the word embedding space of $T_{\theta}$, $\delta_{g}$ is a positive number between 00 and $1$, and $r$ can be any positive number. We then define $\epsilon$ and $\delta$ as below:
ϵ=italic-$\epsilon=$ $\sqrt{d}L\mu+r+\frac{\sqrt{d}\epsilon_{f}}{\mu}$ (9)
$\delta=\delta_{g}+\frac{p_{\theta}(x)(1-p_{\theta}(x))}{\epsilon_{f}^{2}}\frac{1}{N}-\frac{p_{\theta}(x)(1-p_{\theta}(x))}{\epsilon_{f}^{2}}\frac{\delta_{g}}{N},$ (10)
and the approximation error bound of $\|\phi_{\theta}(x)\|$ can be written as:
$\textsf{Pr}(\|g_{\theta}(x)-\nabla\phi_{\theta}(x)\|\leq\epsilon)\geq 1-\delta$
where
[…]
with a proper selection of the smoothing parameter $\mu$. A good approximation error bound is guaranteed by small $\epsilon$ and $\delta$. We can obtain a smaller $\delta_{g}$ and $r$ when $P$ increases according to Equation 8. We then analyze the approximation error as below:
  • When $P$ or $N$ increases, We can decrease $\delta$ by taken smaller value for $\delta_{g}$ or larger value for $N$.
  • When $P$ increases, We can decrease $\epsilon$ by taken smaller value for $r$.
This bound analysis demonstrates that we can reduce the approximation error relative to the true gradient $g_{\theta}(x)$ by taking larger values for $N$ and $P$. ∎
A.11 Ablation study on $P$ and $N$ in Gradient Cuff
Recall that $q$, the total query times to the target LM $T_{\theta}$ in Gradient Cuff, is defined as below:
$q=N\times(P+1)$
We have two strategies to increase $q$:
  • Fixed-N. Keep $N$ fixed and increase $q$ by increasing $P$.
  • Fixed-P. Keep $P$ fixed and increase $q$ by increasing $N$.
Table 5: (N,P) combinations when increasing query times
Strategy Total Query Times 10 20 30 40 fixed-N N=5, P=1 N=5, P=3 N=5, P=5 N=5, P=7 fixed-P N=5, P=1 N=10, P=1 N=15, P=1 N=20, P=1
We set $\sigma=10\%$ to evaluate fixed-N and fixed-P on all types of malicious queries and compare the average refusal rate. In this experiment, we increase $q$ from $10$ to $40$. The exact $(N,P)$ combinations can be found in Table  5.
[…]
Though we can improve Gradient Cuff with fixed-N, the runtime of our algorithm would become much longer due to hundreds of language model calls. We could use batch-inference or other inference-accelerating methods to reduce the runtime. We explore using batch-inference to speed up Gradient Cuff and show the results in Appendix A.12
A.12 Running Time Reduction
Figure 2: Speed up Gradient Cuff with Batch Inference
On LLaMA-7B-Chat, we evaluate the runtime of a defence by applying Gradient Cuff to one query 10 times and averaging the runtime. The length of the query we use is $10$. We choose $N=10$ and $P=10$ so that the total number of queries $q$ for a single pass can be obtained by:
$q=N\times(P+1)=110$
We select the batch size from $1,2,5,10,11,55,110$ to show the improvement in running speed provided by Batch Inference. Figure 2 shows that the time for one run of Gradient Cuff can be greatly reduced by applying batch inference.
A.13 Defending Jailbreak-free Malicious Queries
For completeness, we tested Gradient Cuff on AdvBench and compared it with all the baselines with malicious user instructions w/o jailbreak prompts (that is, naive jailbreak attempts). Specifically, the test set is the collection of the 100 jailbreak templates sampled from AdvBench harmful behaviors. From the results shown in Figure 3, we can see that all methods can attain a good defense performance for those jailbreak-free malicious queries.
Figure 3: Defense performance on jailbreak-free malicious queries. We ran all the experiments under 5 random seeds and reported the average refusal rate.
Algorithm 2 Adaptive PAIR
  Input: Number of iterations $K$, objective $O$, Attacker $A$, Target Model $T$, Judge Function JUDGE
  Initialize the system prompt of $A$ with $O$
  Initialize conversation history $C=[\ ]$
  for $i=1$ to $k$ do
     $P\sim q_{A}(C)$Generate a candidate prompt $P$
[…]
     #Generate the target response $R$
     $S\leftarrow\texttt{JUDGE}(P,R)$Compute the judge score
     if $S$ == JAILBROKEN then
        Return: $P$
     end if
     $C\leftarrow C+[P,R,S]$Add to conversation
  end for
  Return: None
Algorithm 3 Adaptive TAP
  Input: A goal $G$, a branching-factor $b$, a maximum width $w$, and a maximum depth $d$
  Oracles: Query access to an attacker language model $A$, a target language model $T$, and JUDGE and off-topic functions. Initialize the system prompt of $A$ Initialize a tree whose root has an empty conversation history and a prompt $G$
  while depth of the tree is at most $d$ do
     Branch:
     for each leaf $\ell$ of the tree do
        Sample prompts $P_{1},P_{2},\dots,P_{b}\sim q(C;A)$, where $C$ is the conversation history in $\ell$
        Add $b$ children of $\ell$ with prompts $P_{1},\dots,P_{b}$ respectively and conversation histories $C$
     end for
     Prune (Phase 1):
     for each (new) leaf $\ell$ of the tree do
[…]
     end for Query and Assess:
     for each (remaining) leaf $\ell$ of the tree do
[…]
        $P=$ the prompt in node $\ell$
        if $P$ is not rejected by Gradient Cuff then
           Sample response $R\sim q(P;T)$
        end if
        Evaluate score $S\leftarrow\texttt{JUDGE}(R,G)$ and add score to node $\ell$
        If $S$ is JAILBROKEN, then return $P$
        Append $[P,R,S]$ to node $\ell$’s conversation history
     end for
     Prune (Phase 2):
     if the tree has more than $w$ leaves then
        Select the top $w$ leaves by their scores (breaking ties arbitrarily) and delete the rest
     end if
  end while
  Return None
Algorithm 4 Adaptive GCG
  Input: Initial prompt $x_{1:n}$, modifiable subset $\mathcal{I}$, iterations $T$, loss $\mathcal{L}$, $k$, batch size $B$
  Word Embedding Layer of the protected language model $e_{\theta}(\cdot)$, Word Embedding Dimension $d$, Random Gaussian Vector generator $R$, Gradient Cuff perturb number $P$.
  $N=0$
  repeat
     for $i=1$in $\mathcal{I}$ do
        $V=R(P,d)$ generate P vectors each drawn from $\mathcal{N}(\mathbf{0},\mathbf{I})$ with dimension $d$
        $E=[e(x_{1:n}),e(x_{1:n})\oplus V_{1},e(x_{1:n})\oplus V_{2}\ldots,e(x_{1:n})\oplus V_{P}]$
        $\mathcal{X}_{i}:=\mbox{Top-}k(-\nabla_{e_{x_{i}}}\mathcal{L}(E))$Compute top-$k$ promising token substitutions
        if $x_{i}>x_{i+1}$ then
           Swap $x_{i}$ and $x_{i+1}$
        end if
     end for
     for $b=1$ldots $B$ do
        $\tilde{x}_{1:n}^{(b)}:=x_{1:n}$Initialize element of batch
        $\tilde{x}^{(b)}_{i}:=\mbox{Uniform}(\mathcal{X}_{i})$, where $i=\mbox{Uniform}(\mathcal{I})$Select random replacement token
     end for
     $x_{1:n}=\tilde{x}^{(b^{\star})}_{1:n}$, where $b^{\star}=\text{argmin}_{b}\mathcal{L}(\tilde{x}^{(b)}_{1:n})$Compute best replacement
     $N=N+1$
  until $N=T$